Methanol, CH3OH, is produced in the reaction: CO2 + 2H2 → CH3OH. How many moles of methanol can be produced when 4.5 mol of CO reacts with 7.2 mol H2?


1
Expert's answer
2021-10-10T10:27:23-0400

Solution.

First you need to determine which substance is in excess.

"n(CO2) = \frac{4.5}{1} = 4.5 \ mol"

"n(H2) = \frac{7.2}{2} = 3.6 \ mol"

Thus, CO2 is in excess. Therefore

"n(CH3OH) = \frac{n(H2)}{2} = \frac{7.2}{2} = 3.6 \ mol"

Answer:

n(CH3OH) = 3.6 mol